OpenStudy (hesan):

If a β are not real cube roots of unity, then a^5+ β^5+a β =? (a refers to alpha)

OpenStudy (anonymous):

Put: \[\alpha = w\] \[\beta = w^2\]

OpenStudy (hesan):

what's the answer?

OpenStudy (anonymous):

You tell me..

OpenStudy (hesan):

i don't know...

OpenStudy (hesan):

What 'll be the answer after putting α=w β=w2

OpenStudy (anonymous):

What you don't know tell me..

OpenStudy (hesan):

I dont know the answer of this question.. please tell me if u've solved this?

OpenStudy (anonymous):

Firstly put and then show me your work... do this much only..

OpenStudy (hesan):

ok, wait...

OpenStudy (hesan):

I'm getting the answer 0

OpenStudy (anonymous):

You are right if you are getting zero..

OpenStudy (hesan):

If a β are not real cube roots of unity, then we always put α=w and β=w^2?

OpenStudy (anonymous):

Yes..

OpenStudy (anonymous):

You can put also b=w and a = w^2 That does not matter..

OpenStudy (hesan):

okey then what will be the value of (α^4+β^4) + 1/αβ

OpenStudy (anonymous):

You tell me..

OpenStudy (hesan):

i m getting w+w ?

OpenStudy (anonymous):

You are wrong.. it is again zero..

OpenStudy (hesan):

oh yes, i forgot to include 1. yup its zero

OpenStudy (anonymous):

yes it is..

OpenStudy (hesan):

the shortest trick is to put α and β w and w^2 respectively... Thanks!

OpenStudy (anonymous):

FORMULAS TO REMEMBER IN THIS: \[\huge \color{green}{1 + \omega + \omega^2 = 0}\] \[\huge \color{blue} {\omega^3 = 1}\]